OpenStudy (calculusxy):

In the formula A=bh, of b is tripled and h is quadrupled, then A is multiplied by N. Compute N.

OpenStudy (calculusxy):

@phi

OpenStudy (calculusxy):

@SithsAndGiggles

OpenStudy (nikato):

So you have AN=(3b)(4h) Then simplify the right side And then divide both sides by A So now you have N=...

OpenStudy (calculusxy):

How do you simplify the right side?

OpenStudy (calculusxy):

@nikato

OpenStudy (phi):

when you multiply numbers (or letters) you can change the order so, for example, (3b)(4h)= 3*4*b*h most people know how to simplify the 3*4 part

OpenStudy (calculusxy):

So 3*4=12 Now is it 12*b*h?

OpenStudy (phi):

yes, exactly. But in algebra we don't write the multiply signs... just write 12bh (which means 12 times b times h)

OpenStudy (calculusxy):

Okay. Now what do I do?

OpenStudy (phi):

you have AN= 12 bh they want N "by itself" I would divide both sides by A

OpenStudy (calculusxy):

But how do I do that?

OpenStudy (phi):

when you do arithmetic with letters, you just write down "divide by A" for example AN/A = 12bh/A or \[ \frac{AN}{A}= \frac{12bh}{A} \]

OpenStudy (phi):

next you use a very useful idea: *anything divided by itself is 1" in other words A*N/A (also written as A/A * N ) becomes 1*N and 1 times N is N you get N= 12bh/A

OpenStudy (calculusxy):

Thanks :)

OpenStudy (phi):

It would be good to go over that and see if it "sinks in". Math uses these ideas *all the time* and you really should try to learn it.